OpenStudy (anonymous):

compute the limit lim as x approaches 0, where f(0)=1, f'(0)=0, f''(0)=2, g(0)=0, g'(0)=1, and g''(0)=28

OpenStudy (anonymous):

i guess you use l'hopital twice

OpenStudy (anonymous):

scratch that it is wrong

OpenStudy (anonymous):

i know you get 0 and 1/14 but I put 1/14 and it was wrong

OpenStudy (anonymous):

use l'hopital once, get 0

OpenStudy (anonymous):

oh lord scratch that too!!!

OpenStudy (anonymous):

first off we need to know what we are taking the limit of is it \[\lim_{x\to 0}\frac{f(x)}{g(x)}\]??

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

if so the limit does not exist

OpenStudy (anonymous):

ohhhhhh is it because you can get two answers?

OpenStudy (anonymous):

because you do not get \[\frac{0}{0}\] you get \[\frac{1}{0}\] so there is no limit

OpenStudy (anonymous):

no it is because the denominator goes to zero, but the numerator goes to 1

OpenStudy (anonymous):

if you are going to get a finite limit, and the denominator goes to zero, the numerator must go to zero as well

OpenStudy (anonymous):

but isn't that why you go to the second and third derivatives though?

OpenStudy (anonymous):

if the denominator goes to zero but the numerator does not, you cannot go further

OpenStudy (anonymous):

you would use the derivative if your form was \[\frac{0}{0}\] but it is not in this case

OpenStudy (anonymous):

ohh ok got it! THANKS FOR THE HELP!!!

OpenStudy (anonymous):

i don't know how to explain in any other words i am afraid you can only go further if you have 0/0 not if you have 1/0

OpenStudy (anonymous):

oh good. ok yw